OpenStudy (anonymous):

Find the derivative of f(x)=secx/sqrt(2tanx)

OpenStudy (anonymous):

\[f(x)=\frac{\sec(x)}{\sqrt{2\tan(x)}}\]

OpenStudy (anonymous):

yess that's the equation

OpenStudy (anonymous):

good lord your math teacher must hate you i cannot see a simple way to do this you are going to have to use the quotient rule and a raft of algebra

OpenStudy (anonymous):

ahhahha well, the answer is supposed to be \[-\sqrt{2}secx/2\tan ^{2}x\]

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=\sec(x), f'(x)=\sec(x)\tan(x)\] \[g(x)=\sqrt{2\tan(x)},g'(x)=\frac{\sec^2(x)}{\sqrt{2\tan(x)}}\]

OpenStudy (anonymous):

maybe the algebra will work, who knows?

OpenStudy (anonymous):

I tried it like that previously and my answer is still off

OpenStudy (anonymous):

i can't do it wolfram gives a different answer but probably uses some trig identity

OpenStudy (anonymous):

ohh okkk thanks anyway! :)